OpenStudy (anonymous):

graph the linear equation x-3y=6 by finding the intercept

OpenStudy (anonymous):

y=(1/3)x-2

OpenStudy (anonymous):

how did you get that, i dont understand

OpenStudy (lalaly):

x intercept when y=0 (6,0) y intercept when x=0 (0,-2)